OpenStudy (anonymous):

Ms. Thomas is making a badge for her son in the shape of a square. The sides of the badge measure 3 inches each. Ms. Thomas wants to put a stripe across the badge from one vertex to the other vertex. What is the approximate length of the stripe? A. 3.5 inches B. 4.2 inches C. 4.9 inches D. 6.0 inches

OpenStudy (elonasushchik):



OpenStudy (elonasushchik):

you can use Pythagoras to solve for the diagonal (hypotenuse) 3^2 + 3^2 = d^2 ___ 18 = d^2 ___ √18 = d ___ √9 * √2 = d ___ 3 √2 = d B looks like a winner

OpenStudy (elonasushchik):

@kathyS123 did u understand my explanation?!

OpenStudy (anonymous):

yes thank you

OpenStudy (elonasushchik):

Anytime!!!!

OpenStudy (anonymous):

can you help me on one more plz ??

OpenStudy (anonymous):

Farmingdale is located 8 mi due west of Cambridge. Bayville is located due east of Cambridge, and Louisville is located due north of Cambridge. Louisville is 17 mi from Farmingdale and 24 mi from Bayville. How far is Bayville from Cambridge? Round to the nearest tenth of a mile. mi A right triangle with a dot at each corner. The dot at the right angle is labeled Cambridge. The other two dots are labeled Lousville and Bayville. A second triangle shares the dots of Cambridge and Louisville. Its third dot is labeled Farmingdale and forms a line with the Cambridge and Bayville dots.

OpenStudy (anonymous):

i sort of get it i know you have use squire root of A^2 and B^2

OpenStudy (anonymous):

i think it's 18.7 mi i am i right ??